A Cessna 175 can average 130 mph. If a trip takes 2 hours one way and the return takes 1 hour and 15 minutes, find the wind speed, assuming it is constant.
This is an algebra equation. I don't even know where to begin. But I know that the wind speed would be x. I'm not quite sure if I need to add or subtract the two time differences. Thank you so much for helping me.
Melissa

Answer by rfer(16322)

with the wind is x+y
against the wind is x-y
same distance both ways.
2(x+y)=1.25(x-y)
2x+2y=1.25x-1.25y
.75x=3.25y
x=4.33y
--------
2(x+y)=130
replace x with 4.33y
2(4.33y+y)=130
9.66y=130
y=14 mph wind